# Relaxed bound on performance of quantum key repeaters and secure content of generic private and independent bits

### 29 Feb 2024 v3

@@ VII Discussion
We have generalized bound by M. Christandl and R. Ferrara Christandl and Ferrara (2017) to the case of arbitrary key correlated states. We then show a sequence of relaxation of this bound, from which it follows that the repeated key of a key correlated state can not be larger from one-way distillable entanglement than by twice the max-relative entropy of its attacked state.
We further ask how big is the key content of a random private bit, which need not be irreducible Horodecki et al. (2018). A not irreducible private bit can have more distillable key than $1$. It turned out that the amount of key is bounded by a constant factor independent of the dimension of the shielding system. It is interesting if the constant $\approx 0.36$ can be improved. One could also ask if this randomization technique also results in a state for which the repeated key is vanishing with large dimension of the shield as it was shown by a different technique of randomization in Christandl et al. (2020). Recent results on private randomness generation Yang et al. (2019) let us also estimate the private randomness content of generic independent bits. Generalizing this result for independent states of larger dimensions would be the next important step.
We also show a bound on the distillable key based on the techniques developed in this manuscript. The bound can not be better than the bound by $E_{R}$. However, it can possibly report how the relative entropy distance deviates from being a metric. Indeed, our bound is in terms of a ’proxy’ arbitrary state. Indeed, we have two terms: $\frac{\alpha}{1-\alpha}D(\rho||\rho^{\prime})$ and ${\tilde{D}}_{\alpha}(\rho^{\prime}||\sigma)$ where $\rho^{\prime}$ is some intermediate state between $\rho$ and separable state $\sigma$. One could believe that the bound is non-trivial, as the relative entropy functions involved are not metrics. In particular, the triangle inequality reporting that going via proxy state yields a larger result does not hold. However, we know anyway that the bound can not be better than the one by $E_{R}$ itself, hence, conversely, we can perhaps conclude from the bound, how far are the relative entropy functions from being a metric.
Finally, we would like to stress that our approach is generic. That is, it appears that any other strong-converse bound on the quantum distillable key may give rise to a new, possibly tighter bound on the one-way quantum key repeater rate. Hence recent strong-converse bound on quantum privacy amplification Salzmann and Datta (2022) paves the way for future research in this direction. It would be important to extend this technique for the two-way quantum key repeater rate.
@@ Acknowledgements.
KH acknowledges Roberto Ferrara, Siddhartha Das and Marek Winczewski for helpful discussion. KH acknowledges the Fulbright Programm and Mark Wilde for hospitality during the Fulbright scholarship at the School of Electric and Computer Engineering of the Cornell University. We acknowledge Sonata Bis 5 grant (grant number: 2015/18/E/ST2/00327) from the National Science Center. We acknowledge partial support by the Foundation for Polish Science (IRAP project, ICTQT, contract no. MAB/2018/5, co-financed by EU within Smart Growth Operational Programme). The ’International Centre for Theory of Quantum Technologies’ project (contract no. MAB/2018/5) is carried out within the International Research Agendas Programme of the Foundation for Polish Science co-financed by the European Union from the funds of the Smart Growth Operational Programme, axis IV: Increasing the research potential (Measure 4.3).
Appendix A Bound for distillable key
We derive now a novel form of a bound on the distillable key in terms of relative entropy functions. Although the bound reduces to a well-known bound $K_{D}\leq E_{R}$ Horodecki et al. (2005, 2009a), we state it due to its novelty. The novelty comes from the fact that it involves the relative entropy distance from arbitrary, not necessarily separable, quantum state. The main result is encapsulated in Theorem 86 below.
We first focus on the analogue of Lemma 1 and Lemma 2 for the key distillation rather than repeated key distillation. The abstract version of the Lemma 1 reads
Lemma 3.
For a bipartite state of the form $\tilde{\sigma}:=\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\sigma_{ABA^{\prime}B^{\prime}}\tau^{\dagger}$, where $\sigma_{ABA^{\prime}B^{\prime}}$ is arbitrary state and $\tau$ is a twisting, there is
$\operatorname{Tr}\tilde{\sigma}\psi_{+}^{\otimes m}\leq 2^{-(\frac{\alpha-1}{\alpha})[m-\tilde{E}_{\alpha}(\sigma)]},$ (79)
where $m\geq 1$ is a natural number, $\psi_{+}$ denotes the maximally entangled state, and $\alpha\in(1,\infty)$.
Proof.
As in the proof of the Lemma 1, we directly use the proof of the strong-converse bound for the distillable key of Das et al. (2021) which is based on Das et al. (2020); Christandl and Müller-Hermes (2017). Let us suppose $\operatorname{Tr}\tilde{\sigma}\psi_{+}^{\otimes m}=1-\epsilon$. Let us also choose a state $\hat{\sigma}$, to be a twisted separable state of the form. $\hat{\sigma}=\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\sigma_{AA^{\prime}:BB^{\prime}}^{\prime}\tau^{\dagger}$ with $\sigma^{\prime}$ being arbitrary separable state in $(AA^{\prime}:BB^{\prime})$ cut. Since any such state has overlap with the singlet state $\psi^{\otimes m}_{+}$ less than $\frac{1}{2^{m}}\equiv 1/K$ (see lemma 7 of Horodecki et al. (2009a)), we have:
$D^{\epsilon}_{h}(\tilde{\sigma}||\hat{\sigma})\geq\log_{2}K=m$ (80)
and further for any $\alpha\in(1,\infty)$
$\tilde{D}_{\alpha}(\tilde{\sigma}||\hat{\sigma})+\frac{\alpha}{\alpha-1}\log_{2}(\frac{1}{1-\epsilon})=$
$\tilde{D}_{\alpha}(\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\sigma\tau^{\dagger}||\operatorname{Tr}_{A^{\prime}B^{\prime}}U\sigma^{\prime}_{A:B}\tau^{\dagger})+$
$\frac{\alpha}{\alpha-1}\log_{2}(\frac{1}{1-\epsilon})\geq m$ (81)
[…]
$\tilde{D}_{\alpha}(\sigma||\sigma^{\prime})+\frac{\alpha}{\alpha-1}\log_{2}(\frac{1}{1-\epsilon})\geq m$ (82)
Since $\sigma^{\prime}$ is an arbitrary separable state, we can take also the infimum over the set of separable states obtaining:
$\tilde{E}_{\alpha}(\sigma)+\frac{\alpha}{\alpha-1}\log_{2}(\frac{1}{1-\epsilon})\geq m.$ (83)
We can now rewrite it as follows:
$1-\epsilon\leq 2^{-(\frac{\alpha-1}{\alpha})(m-E_{\alpha}(\sigma))}$ (84)
Since $1-\epsilon=\operatorname{Tr}\tilde{\sigma}\psi_{+}^{\otimes m}$, the assertion follows. ∎
Lemma 4.
For a maximally entangled state $\psi_{+}^{\otimes m}$, a state $\tilde{\sigma}:=\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\sigma\tau^{\dagger}$ where $\tau$ is a twisting, for every $\alpha\in(1,\infty)$, there is
[…]
Proof.
The proofs goes in full analogy to the proof of Lemma 2. ∎
Theorem 7.
For every bipartite state $\rho\in{\cal B}({\cal C}^{d_{A}}\otimes{\cal C}^{d_{B}})\equiv{\cal H}$, there is:
$K_{D}(\rho)\leq\inf_{\sigma\in{\cal H},\alpha\in(1,\infty]}[\frac{\alpha}{\alpha-1}D(\rho||\sigma)+\tilde{E}_{\alpha}^{\infty}(\sigma)]$ (86)
Proof.
The proof goes along the line of the proof of the fact that $E_{R}$ is the upper bound on distillable key $K_{D}$ Horodecki et al. (2005, 2009a). Let the LOCC protocol $\Lambda$ distill a state close by $\epsilon>0$ to a private state $\gamma_{m}$ with $m$-qubit key part, when acting on $n$ copies of $\rho$. Let us also denote $m=nR$ where $R>K_{D}(\rho)-\eta$ by some $\eta>0$. We will show the chain of (in)equalities and comment them below:
$D(\rho||\sigma)={1\over n}D(\rho^{\otimes n}||\sigma^{\otimes n})\geq$ (87)
${1\over n}D(\Lambda(\rho^{\otimes})||\Lambda(\sigma^{\otimes n}))\geq$ (88)
${1\over n}D(\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\Lambda(\rho^{\otimes})\tau^{\dagger}||\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\Lambda(\sigma^{\otimes n})\tau^{\dagger})\equiv$
${1\over n}D(\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\Lambda(\rho^{\otimes})\tau^{\dagger}||\sigma_{n})\geq$ (89)
${1\over n}\inf_{\tilde{\sigma}_{n}\in Cone(SEP,\sigma_{n})}D(\tilde{\psi}_{+}^{\otimes m}||\tilde{\sigma}_{n})\geq$ (90)
$\inf_{\tilde{\sigma}_{n}\in Cone(SEP,\sigma_{n})}D(\psi_{+}^{\otimes m}||\tilde{\sigma}_{n})+{1\over n}O(\epsilon+h(\epsilon))\geq$ (91)
${1\over n}[n\frac{\alpha-1}{\alpha}(R-{1\over n}\tilde{E}_{\alpha}(\sigma^{\otimes n})]+{1\over n}O(\epsilon+h(\epsilon))$ (92)
In the first step, we use the tensor property of the relative entropy. In the next, we use the monotonicity of the relative entropy under $\Lambda$, and further under a map which consists of the inverse to the operation twisting by $\tau$ and tracing out system $A^{\prime}B^{\prime}$. The twisting $\tau$ is defined by $\gamma_{m}$. We further relax relative entropy to infimum over states from $Cone(SEP,\sigma_{n})$ to which $\sigma_{n}$ belongs, and use asymptotic continuity of the relative entropy from a convex set. We then use the analogue of Lemma 2.
Taking the limit of $n$ going to infinity, $\eta\rightarrow 0$ and $\epsilon\rightarrow 0$ proves the thesis. ∎
Remark 1.
The bound of Eq. (86) has two extreme, however equivalent cases. First is when we set $\sigma=\rho$. In that case the first term equals zero because $D(\rho||\rho)=0$, and the second term equals $\tilde{E}_{\alpha}(\rho)$, where $\alpha>1$. Taking limit $\alpha\rightarrow 1$, we obtain a known bound $E_{R}(\rho)$ on distillable key. The second extreme is when we set $\sigma$ to be any separable state. Then the second term is zero as $\tilde{E}_{\alpha}$ is the relative entropy divergence ”distance” from the set of separable states. In that case, we can take the limit $\alpha\rightarrow\infty$, and again obtain the bound $E_{R}(\rho)$.
Appendix B Random quantum objects
@@
In this section we provide a short introduction into random matrices. The scope is limited to concepts necessary in the understanding of our result.
B.1 Ginibre matrices
@@
We start of by introducing the Ginibre random matrices ensemble Ginibre (1965). This ensemble is at the core of a vast majority of algorithms for generating random matrices presented in later subsections. Let $(G_{ij})_{1\leq i\leq m,1\leq j\leq n}$ be a $m\times n$ table of independent identically distributed (i.i.d.) random variable on $\mathbb{C}$. The field $\mathbb{K}$ can be either of $\mathbb{R}$, $\mathbb{C}$ or $\mathbb{Q}$. With each of the fields we associate a Dyson index $\beta$ equal to $1$, $2$, or $4$ respectively. Let $G_{ij}$ be i.i.d random variables with the real and imaginary parts sampled independently from the distribution $\mathcal{N}(0,\frac{1}{\beta})$. Hence, $G\in\mathrm{L}(\mathcal{X},\mathcal{Y})$, where matrix $G$ is
@@
B.2 Wishart matrices
Wishart matrices form an ensemble of random positive semidefinite matrices. They are parametrized by two factors. First is the Dyson index $\beta$ which is equal to one for real matrices, two for complex matrices and four for symplectic matrices. The second parameter, $K$, is responsible for the rank of the matrices. They are sampled as follows
@@ B.2 Wishart matrices
B.3 Circular unitary ensemble
@@
Another approach is to consider a Ginibre matrix $G\in\mathrm{L}(\mathcal{X})$ and its polar decomposition $G=UP$, where $U\in\mathrm{L}(\mathcal{X})$ is unitary and $P$ is a positive matrix. The matrix $P$ is unique and given by $\sqrt{G^{\dagger}G}$. Hence, assuming $P$ is invertible, we could recover $U$ as
@@
The optimal approach is to utilize the QR decomposition of $G$, $G=QR$, where $Q\in\mathrm{L}(\mathcal{X})$ is unitary and $R\in\mathrm{L}(\mathcal{X})$ is upper triangular. This procedure is unique if $G$ is invertible and we require the diagonal elements of $R$ to be positive. As typical implementations of the QR algorithm do not consider this restriction, we must enforce it ourselves. The algorithm is as follows
@@
B.4 Random mixed quantum states
@@
Random mixed states can be generated in one of two equivalent ways. The first one comes from the partial trace of random pure states. Suppose we have a pure state |$\ket{\psi}\in\mathcal{X}\otimes\mathcal{Y}$. Then we can obtain a random mixed as
@@
It can be easily verified that this approach is equivalent to the one utilizing random pure states. First, note that in both cases, we start with $\dim(\mathcal{X})\dim(\mathcal{Y})$ complex random numbers sampled from the standard normal distribution. Next, we only need to note that taking the partial trace of a pure state |$\ket{\psi}$ is equivalent to calculating $AA^{\dagger}$ where $A$ is a matrix obtained from reshaping |$\ket{\psi}$.
The properties of these states have been extensively studied. We will omit stating all the properties here and refer the reader to Wootters (1990); Życzkowski and Sommers (2001); Sommers and Życzkowski (2004); Puchała et al. (2016); Zhang et al. (2017); Zhang (2017).

### 09 Mar 2026 v4

@@ VII Discussion
We have generalized the bound by M. Christandl and R. Ferrara Christandl and Ferrara (2017) to the case of arbitrary key correlated states. We then show a sequence of relaxation of this bound, from which it follows that the repeated key of a key correlated state can not be larger than twice one-way distillable entanglement plus the max-relative entropy of its attacked state.
We further ask how big the key content of a random private bit is, which need not be irreducible Horodecki et al. (2018). A not irreducible private bit can have more distillable key than $1$. It turned out that the amount of key is bounded by a constant factor independent of the dimension of the shielding system. It is interesting if the constant $\approx 0.36$ can be improved. One could also ask if this randomization technique also results in a state for which the repeated key is vanishing with a large dimension of the shield, as it was shown by a different technique of randomization in Christandl et al. (2020). Recent results on private randomness generation Yang et al. (2019) also allow us to estimate the private randomness content of generic independent bits. Generalizing this result to independent states of larger dimensions would be the next important step. Finally, we emphasize that our approach is generic. That is, it appears that any other strong-converse bound on the quantum distillable key may give rise to a new, possibly tighter bound on the one-way quantum key repeater rate. Hence, recent strong-converse bound on quantum privacy amplification Salzmann and Datta (2022) paves the way for future research in this direction. It would be important to extend this technique to the two-way quantum key repeater rate.
@@ Acknowledgements.
KH acknowledges Roberto Ferrara, Siddhartha Das and Marek Winczewski for helpful discussion. KH acknowledges the Fulbright Programm and Mark Wilde for hospitality during the Fulbright scholarship at the School of Electric and Computer Engineering of the Cornell University. We acknowledge Sonata Bis 5 grant (grant number: 2015/18/E/ST2/00327) from the National Science Center. We acknowledge partial support by the Foundation for Polish Science (IRAP project, ICTQT, contract no. MAB/2018/5, co-financed by EU within Smart Growth Operational Programme). The ’International Centre for Theory of Quantum Technologies’ project (contract no. MAB/2018/5) is carried out within the International Research Agendas Programme of the Foundation for Polish Science co-financed by the European Union from the funds of the Smart Growth Operational Programme, axis IV: Increasing the research potential (Measure 4.3). KH acknowledges National Science Centre, Poland grant OPUS UMO-2023/49/B/ST2/02468.
Appendix A Random quantum objects
@@
In this section, we provide a short introduction to random matrices. The scope is limited to concepts necessary for the understanding of our result.
A.1 Ginibre matrices
@@
We start off by introducing the Ginibre random matrices ensemble Ginibre (1965). This ensemble is at the core of a vast majority of algorithms for generating random matrices presented in later subsections. Let $(G_{ij})_{1\leq i\leq m,1\leq j\leq n}$ be a $m\times n$ table of independent identically distributed (i.i.d.) random variables on $\mathbb{C}$. The field $\mathbb{K}$ can be either of $\mathbb{R}$, $\mathbb{C}$ or $\mathbb{Q}$. With each of the fields, we associate a Dyson index $\beta$ equal to $1$, 22, or $4$, respectively. Let $G_{ij}$ be i.i.d random variables with the real and imaginary parts sampled independently from the distribution $\mathcal{N}(0,\frac{1}{\beta})$. Hence, $G\in\mathrm{L}(\mathcal{X},\mathcal{Y})$, where matrix $G$ is
@@
A.2 Wishart matrices
Wishart matrices form an ensemble of random positive semidefinite matrices. They are parametrized by two factors. First is the Dyson index $\beta$, which is equal to one for real matrices, two for complex matrices, and four for symplectic matrices. The second parameter, $K$, is responsible for the rank of the matrices. They are sampled as follows
@@ A.2 Wishart matrices
A.3 Circular unitary ensemble
@@
Another approach is to consider a Ginibre matrix $G\in\mathrm{L}(\mathcal{X})$ and its polar decomposition $G=UP$, where $U\in\mathrm{L}(\mathcal{X})$ is unitary, and $P$ is a positive matrix. The matrix $P$ is unique and given by $\sqrt{G^{\dagger}G}$. Hence, assuming $P$ is invertible, we could recover $U$ as
@@
The optimal approach is to utilize the QR decomposition of $G$, $G=QR$, where $Q\in\mathrm{L}(\mathcal{X})$ is unitary and $R\in\mathrm{L}(\mathcal{X})$ is upper triangular. This procedure is unique if $G$ is invertible, and we require the diagonal elements of $R$ to be positive. As typical implementations of the QR algorithm do not consider this restriction, we must enforce it ourselves. The algorithm is as follows
@@
A.4 Random mixed quantum states
@@
[…]
@@
[…]
The properties of these states have been extensively studied. We will omit stating all the properties here, and refer the reader to Wootters (1990); Życzkowski and Sommers (2001); Sommers and Życzkowski (2004); Puchała et al. (2016); Zhang et al. (2017); Zhang (2017).